OpenStudy (anonymous):

Verify the identity sec^2 x cot x - cot x = tan x

OpenStudy (lgbasallote):

by verify you mean to prove right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

well i'll go with the left hand side since it is more complicated..you'd like to do that part in proving.. so...factor out cotx first.. \[\large \mathbf{cotx(sec^2x - 1)}\] do you know what to do next? \(\large \mathsf{\text{Hint:}}\) \(\large \mathsf{tan^2x + 1 = sec^2x}\)

OpenStudy (anonymous):

make sec^2x-1 tan^2x?

OpenStudy (anonymous):

not minus but exchange

OpenStudy (anonymous):

cotx(tan^2 x)

OpenStudy (lgbasallote):

right! now change to sines and cosines...

OpenStudy (anonymous):

i have (cos x/sin x)(1/cos^2 x)

OpenStudy (lgbasallote):

you got the cotx right...but tan^2 x is NOT 1/cos^2 x...try again

OpenStudy (anonymous):

woopsies sec^2x -1 is tan^2x

OpenStudy (lgbasallote):

no no..change tan^2 x in terms of sines and cosines..

OpenStudy (anonymous):

which is \[\sin ^{2} x/\cos ^{2} x\]

OpenStudy (lgbasallote):

right! so... \(\Huge (\frac{cos}{sin})(\frac{sin^{2}x}{cos^{2}x})\) what do you think comes next?

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

cancel out

OpenStudy (anonymous):

leaving only sin x/cos x

OpenStudy (anonymous):

which is tan x

OpenStudy (lgbasallote):

\[\Huge \checkmark\]

OpenStudy (anonymous):

yaaaaay

OpenStudy (anonymous):

thank you for making me try, that helped out instead of just telling me the answer!!

OpenStudy (lgbasallote):

well that is what all users of OpenStudy must do :DDD <tips hat>

OpenStudy (anonymous):

ok, on to solving trig equations dun dun dun

OpenStudy (lgbasallote):

\(\large \mathsf{\text{Goodluck! Have Fun!}}\) :DDD

OpenStudy (anonymous):

:)

OpenStudy (lgbasallote):

...i would appreciate some medal for the hard work *whistles* lol haha =)) you should award people for their hard work yknow ;D

OpenStudy (anonymous):

i just started this website today! I have no idea how to use it yet

OpenStudy (anonymous):

how do i give you a medal

OpenStudy (anonymous):

ill give you 1000

OpenStudy (anonymous):

lol

OpenStudy (lgbasallote):

ha thanks :D and by the way... \(\Huge \mathbb{\text{WELCOME TO OPENSTUDY!!!}}\) remember to give medals to those you think gave you the best answer/explanation ;D

OpenStudy (lgbasallote):

^hope you can rad that..coz it got cut in mine

OpenStudy (lgbasallote):

read*

OpenStudy (anonymous):

i love it! but how do you give medals? by clicking best answer?

OpenStudy (lgbasallote):

yepp :DD

OpenStudy (lgbasallote):

you might wanna check this site http://openstudy.com/study#/groups/OpenStudy%20Feedback read some of the posts of the users with the violet names (they're the site moderators) you can learn about the current functions in OpenStudy from those posts :DD

OpenStudy (anonymous):

i wish i knew about open study a year ago

OpenStudy (lgbasallote):

haha lol yeah..this is a great site to learn ^_^